Consider the following LP problem.

MAX: 9X1-8X2

Subject to: x1+x2≤6

-x1+x2≤3

3x1-6x2≤4

x1,x2≥0

Sketch the feasible region for this model.

What is the optimal solution?

What is the optimal solution if the objective function changes to Max.-9x1+8x2?
Answer #1

Replace inequality with equal sign in the constraints, solve the equations by giving one variable a hypothetical numerical value.

x1

x2

x1+x2=6

-1

7

6

0

-x1+x2=3

3

6

-6

-3

3x1-6x2=4

-6

-2.3

5.3

2

Please note that the 4th constraint of x1,x2>=0, gives us the Y axis as a constraint to consider.

00 -1,7 3,6 (4.4,1.6) Feasible Region (3,0) 15.3,2 (1.8,4.4) 16,0 0 (0,0) 61-6,-2.3 -6,-3

hence, the 4 options for x1 and x2 are:

x1

x2

9x1-8x2

4.4

1.6

26.8

1.8

4.4

-19

0

0

0

3

0

27

since the objective is to maximize the objective function, the maximum value is given by (4.4,1.6), hence, the value of x1=4.4, x2=1.6

Optimal solution

x1=

4.4

x2=

1.6

if the objective function is -9x1+8x2

x1

x2

9x1-8x2

4.4

1.6

-26.8

1.8

4.4

19

0

0

0

3

0

-27

since the objective is to maximize the objective function, the maximum value is given by (1.8,4.4),

Optimal solution

x1=

1.8

x2=

4.4

1 Replace inequality with equal sign in the constraints, solve the equations by giving one variable a hypothetical numerical